OpenStudy (anonymous):

Integral of tan^4 (x) * sec^2 (x) dx

OpenStudy (anonymous):

derivative of tan(x) is ?

OpenStudy (anonymous):

sec^2 (x)

OpenStudy (anonymous):

yes, so your u-substitution is u= ?

OpenStudy (welshfella):

- thats the way to go

OpenStudy (anonymous):

Ohh, how do we determine the u substitution?

OpenStudy (anonymous):

you know that for any \[\int\limits_{ }^{ } F\left(~g(x)~\right)~~\times ~g'(x)~~dx\] then you set: u=g(x) du=g'(x) dx and you get: \[\int\limits_{ }^{ } F\left(~u~\right) ~du\]

OpenStudy (anonymous):

is this last post making sense?

OpenStudy (anonymous):

YEsssss

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

tan^4(x) is same as (tan x)^4

OpenStudy (anonymous):

so your u= ?

OpenStudy (anonymous):

tanx

OpenStudy (anonymous):

yes, and your new integral is ?

OpenStudy (anonymous):

integral of tan^4(x) dx

OpenStudy (anonymous):

Yes I understand now, thank you

OpenStudy (anonymous):

no, not really...

OpenStudy (anonymous):

\[\int\limits_{ }^{ } (\color{red}{\tan x})^4\cdot \color{green}{(\sec^2x)~dx}\]\[\int\limits_{ }^{ } (\color{red}{u})^4 \color{green}{du}\]

OpenStudy (anonymous):

the red is our u-substitution and green is what "du" is